# Diving into Self-Evolving Training for Multimodal Reasoning

### v1

@@ Abstract
Reasoning ability is essential for Large Multimodal Models (LMMs). In the absence of multimodal chain-of-thought annotated data, self-evolving training, where the model learns from its own outputs, has emerged as an effective and scalable approach for enhancing reasoning abilities. Despite its growing usage, a comprehensive understanding of self-evolving training, particularly in the context of multimodal reasoning, remains limited. In this paper, we delve into the intricacies of self-evolving training for multimodal reasoning, pinpointing three key factors: Training Method, Reward Model, and Prompt Variation. We systematically examine each factor and explore how various configurations affect the training’s effectiveness. Our analysis leads to a set of best practices for each factor, aimed at optimizing multimodal reasoning. Furthermore, we explore the Self-Evolution Dynamics during training and the impact of automatic balancing mechanisms in boosting performance. After all the investigations, we present a final recipe for self-evolving training in multimodal reasoning, encapsulating these design choices into a framework we call M-STaR (Multimodal Self-evolving Training for Reasoning), which is universally effective for models with different sizes on various benchmarks, e.g., surpassing the pre-evolved model significantly on 5 multimodal reasoning benchmarks without using additional human annotations, as demonstrated on MiniCPM-V-2.5 (8B), Phi-3.5-Vision (4B) and InternVL2 (2B). We believe this study fills a significant gap in the understanding of self-evolving training for multimodal reasoning and offers a robust framework for future research. Our policy and reward models, as well as the collected data, is released to facilitate further investigation in multimodal reasoning.111https://github.com/hkust-nlp/mstar
Project Page: https://mstar-lmm.github.io
@@ 1 Introduction
With the rapid advancement of Large Language Models, their reasoning abilities have improved significantly (Shao et al., 2024; Xin et al., 2024; Yang et al., 2024). This progress has been accompanied by a growing demand for more realistic and general reasoning capabilities. Multimodal reasoning, considered a fundamental skill in many real-world applications, such as intelligent agents (Liu et al., 2024c), robotics (Li et al., 2023; Liu et al., 2024b), and autonomous driving (Yang et al., 2023), exemplifies this trend. Multimodal reasoning requires Large Multimodal Models (LMMs) to understand various modalities beyond text. For example, visual mathematical reasoning (Lu et al., 2023) challenges models to analyze complex figures, diagrams, and charts, leveraging the provided information to perform reasoning tasks.
@@
Despite these advances, the availability of human-annotated thought processes in multimodal scenarios remains limited, challenging the learning of multimodal reasoning. Consequently, self-evolving training, which utilizes model’s own generation ability to iteratively tune and improve itself without external annotated data, has emerged as an appealing candidate to facilitate reasoning abilities. While research on self-evolving training has primarily focused on the text-only settings (Hosseini et al., 2024; Sun et al., 2024; Shao et al., 2024), its application in the multimodal domain, especially for reasoning tasks, has been limited with only a few sporadic examples (Fang et al., 2024; Dubey et al., 2024; Deng et al., 2024), and a unified framework has yet to be established.
@@
Inspired by the fact that self-evolving training can actually be formulated as a general reinforcement learning (RL) framework (§ 2), we first identify the three key components of self-evolving training in multimodal reasoning via the lens of RL, namely the training method, the use of reward model and the prompt variation, to build a clear and unified design space for searching the optimal strategies. Through massive controlled experiments, we (1) propose a continuous self-evolving training scheme that inherits optimizers states from the last iteration to reduce the gap towards full online learning and outperforms other iterative baselines (§3.2); (2) train the first multimodal, process-based reward model for multimodal reasoning and demonstrate its usefulness in further enhancing performance (§3.3); and (3) find that adding more unlabeled queries helps only when having perfect reward signals (e.g., the oracle groundtruth answers), and it hurts the performance if the reward model does not generalize well on unseen data (§3.4).
Orthogonal to the static factors above, we also dive into the dynamic evolving process of the model, namely the dynamics of self-evolution to see how model’s behavior changes during the training. We find that although the model performance of greedy decoding increases, its exploration potential keeps decreasing through the training (§4.1), then we introduce a mechanism to dynamically adjust the sampling temperature during training to balance exploration and exploitation (§4.2), as shown in Figure 1.
@@
Combining all the recipes concluded through separate, controlled studies, we propose our self-evolving training algorithm named as M-STaR (Multimodal Self-evolving Training for Reasoning). Our experimental results on 5 different multimodal reasoning benchmarks, including MathVista, M3CoT, MMStar, MMBench and AI2D, show that this strategy, which incorporates both optimized static design choices and dynamic adjustments, effectively mitigates exploration loss during training and further enhances performance universally for models with different sizes such as MiniCPM-V-2.5 (8B), Phi-3.5-Vision (4B) and InternVL2 (2B).
@@ 2 Overview of Self-Evolving Training for Multimodal Reasoning
where the ground-truth answer input to the reward function can be empty, for example, when dealing with unlabeled inputs, and then a reward model will be necessary to score .
The Design Spaces
There are different design choices to model and implement Eq. 2, for example, the design of reward function and whether to incorporate additional unlabeled inputs without into training. Additionally, the training algorithms to perform this iterative process vary as well. For example, while Gulcehre et al. (2023); Xu et al. (2024b) initialize the model from the last checkpoint at each iteration, Zelikman et al. (2022); Singh et al. (2023) argue that initializing from the beginning checkpoint reduces overfitting and gives better performance empirically. Next, we investigate these three design spaces, training method, reward model, and prompt variation, aiming to summarize the best practices for each factor to faciliate multimodal reasoning learning.
@@ 3.1 General Setup
Model
We base our main exploration on MiniCPM-V-2.5 (8B) (Yao et al., 2024), and we also validate the final design choice for each component on two extra models with different sizes: Phi-3.5-Vision (4B) (Abdin et al., 2024) and InternVL-2 (2B) (Chen et al., 2024c). The details of these models can be found in Appendix A. To make the analysis process easier to understand, we mainly present the results of MiniCPM-V-2.5 in this section, while we include the results of the other models in §4.2.
Datasets
We utilize MathV360K (Shi et al., 2024), a high-quality and diverse multimodal reasoning dataset as our seed training dataset. Specifically, we downsample half of the examples (180K) from it to serve as our labeled training set, while setting aside the remaining half as a unlabeled training set by not using the answers in it. For evaluation, we split 750 samples from the unlabeled part of MathV360K as the in-domain (ID) testset. For our out-of-domain (OOD) testset we use the testmini split of MathVista (Lu et al., 2023), a widely recognized comprehensive benchmark encompassing a wide range of multimodal reasoning tasks, including visual question answering, figure-based question answering, science question answering, and more. We also keep an non-overlapping 250 samples from MathV360K as the global validation set in training.
Warm-Up Phase to Unlock the Chain-of-Thought (CoT) Capability of LMMs
In our preliminary experiments, we found that open-source LMMs would directly output the answer given the query, while struggling to produce detailed chain-of-thought (CoT) reasoning processes. This may originate from the the scarcity of high quality rationales in most existing multimodal SFT training datasets (Masry et al., 2022; Shi et al., 2024), which limits the ability of open-source LMMs to generate detailed, step-by-step reasoning. Self-evolving training, however, requires responses with varying intermediate steps to allow models to learn effectively from on-policy data. To address this issue, we initiate a warm-up phase as the first step before self-evolving training. Instead of prompting the model to answer questions directly, we prompt it to generate intermediate reasoning steps for a given triplet (question, image, and answer). For each triplet, we ask models to rollout 16 samples with . We then filter out results where the final answers do not match the ground truth and sample 100K from the generated dataset to create a warm-up CoT dataset with correct answers. Finally, we fine-tune our models on this dataset, treating it as a standard RFT process. Our iterative self-evolving training process will then start from this model checkpoint after the warm-up training. For the prompt during the warm-up phase, please refer to Appendix B for more details.
Main Training Settings
We adopt most of the training settings from Yao et al. (2024) (see Appendix C), using a constant learning rate of and training for 10K steps across all experiments. During all rollout phases in training, we sample 16 responses per query and set the sampling temperature to 1.0. Unless explicitly stated otherwise, we follow existing practices (Singh et al., 2023; Zelikman et al., 2022) and only use the labeled training data.
@@ 3.2 Training Methods
Setup
We perform controlled experiments to study the effect of different training methods, thus in this experiment we use the labeled dataset only and simply adopt the binary exact-match reward between ground-truth answer and the generated answer. We compare with the most common iterative self-evolving algorithms ReSTEM (Singh et al., 2023) and iterative RFT, which are specific instantiations of our training methods design space. To study the effect of iteration interval in the proposed continuous self-evolving, we experiment with different percentage of all the queries per iteration, varying from [6.25%, 12.5%, 25%, 50%, 100%].
| Method | Iteration Interval (#) | MathV360K | MathVista | ||
@@
| Iterative RFT | 100%(180K) | 42.3 | 55.7 | ||
| RestEM | 100%(180K) | 42.3 | 55.1 | ||
| Continous Self-Evolving | 100%(180K) | 42.2 | 56.7 | ||
| 50%(90K) | 43.1 | 56.2 | |||
| 25%(45K) | 43.1 | 57.2 | |||
| 12.5%(22K) | 42.3 | 56.1 | |||
| 6.25%(11K) | 41.0 | 56.8 |
Results
Table 1 presents the experimental results of various training methods. Overall, initializing training from the last policy model checkpoint and maintaining a continuous optimization process contribute most significantly to the effectiveness of self-evolving training, particularly on MathVista. Continuous self-evolving achieves the best performance both on the in-domain MathV360K test set, with 43.1%, and on the OOD test set, MathVista, with 57.2%. We also see the importance of maintaining a proper interval to traverse the data queries. With a large interval, the training method becomes closer to an offline one, and the model cannot get timely updates on data matching its current output distribution. On the other hand, switching over the Improve and Generate steps too frequently makes the learning process unstable, leading to a lower score, especially on the in-domain test set. The strategy of continuous self-evolving with proper intervals also works for other smaller models, as shown in Table 4 compared with representative baselines, indicating its effectiveness and generalizability across different model sizes.
@@ 3.3 Reward Models
Setup
We conduct controlled experiments to assess the impact of incorporating the Process Reward Model (PRM) into self-evolving training and explore how best to utilize the reward signals provided by PRM. Notably, before applying PRM, responses are pre-filtered based on their final answers to ensure consistency and quality during training. To train our PRM, we use Monte Carlo rollouts starting from prefixes with partial reasoning steps (Wang et al., 2024) to generate the training data. Specifically, we sample 16 responses per question and complete each step 8 times to obtain step-level annotations. For additional details on the training process of our PRM, please refer to Appendix D. We evaluate two different operations: (1) Top-K: Pick the top-K correct responses according to their reward scores, and (2) Filtering by a Threshold : Filtering out sampled responses with lower aggregated rewards than . The optimal value of is which is determined by enumerating it with an interval of 0.1 on the validation set. Additionally, we investigate how varying the value of in affects training, as it represents a trade-off between the quality and diversity of the samples. According to §3.2, we fix training methods as continuous self-evolving with 45k interval and set continuous self-evolving, with or without randomly selected correct responses as our baselines.
Results
Table 2 presents the results of integrating the PRM into self-evolving training, along with the impact of different choices. Continuous Self-Evolving with PRM using achieves the best performance in both the ID and OOD tests, with scores of 45.3% and 59.2%, respectively. Compared to training without PRM, most instances of self-evolving training with PRM show improved performance, especially in the OOD test. Interestingly, randomly selecting a subset of correct responses actually leads to worse performance than continuous self-evolving, suggesting that even correct answers can be noisy. Random selection may increase the proportion of these noisy samples, undermining the effectiveness of self-evolving training.
In terms of leveraging PRM, we found that using with a moderate number of responses—a re-ranking operation to select correct responses with the highest-quality intermediate steps—outperforms filtering by a threshold. The results also highlight the importance of balancing the quality and diversity of sampled responses in self-evolving training. Selecting strikes this balance well, ensuring both response diversity and high-quality reasoning steps for each question. Similar to the results in §3.2, we also see improvement when involving PRM on smaller models in Table 4.
@@ Results
| Continuous Self-Evolving | - | 43.1 | 57.2 | |
| + Random Selection | 41.0 | 55.5 | ||
@@
What makes PRM work for self-evolving training?
To pursue deeper insights into the role of PRM in self-evolving training, we conduct an analysis presented in Figure 2. Based on the experimental results from §3.3, we explore PRM’s impact from two key perspectives: (1) Can PRM help the model to select out correct responses among different numbers of rollouts? (2) How different are the Top 2 and the rest correct solutions re-ranked by reward scores? We use the first checkpoint after warmup as policy model to sample 16 responses for each question in the validation set with temperature= and reveal the behaviors of PRM in these samples.
@@
To understand why our PRM can still significantly contributes to self-evolving training despite its weaker verification abilities, we analyzed the distribution of other metrics for the top-2 selected responses compared to other correct responses. We approached this from two perspectives: the average number of reasoning steps, and how much a response is directly relevant to the question annotated by GPT-4o (see Appendix E), since we do not find incorrect steps but find some irrelevant steps after randomly checking some examples . As shown in Figures 2(b) and 2(c), the responses re-ranked by our PRM generally have fewer reasoning steps and more relavant to the query. This highlights the precision of our PRM in recognizing genuinely high-quality responses. Therefore, our PRM acts as an effective reranker, precisely identifying top-quality responses. This precision is especially critical in self-evolving training, where responses are already filtered by ground-truth answers, and the ability to accurately assess the quality of reasoning steps becomes vital.
@@ 3.4 Prompt Variation
Skylines: Unlabeled Prompts with Oracle Reward Signals
The coupling of these additional factors introduces complexity, making the effective use of unlabeled prompts less predictable. To dissect these factors, we start by establishing a baseline with “skyline” experiments, where both the unlabeled prompts and their ground truth answers are available but not used during the SFT phase. These unlabeled prompts with oracle reward signals serve as an intermediate difficulty between fully unlabeled and labeled prompts, providing insight into the challenges of training with unlabeled data.
Unlabeled Prompts
We incorporate unlabeled prompts into self-evolving training. To ensure the quality of sampled responses for these prompts, we use weighted voting to ensemble the predictions from different responses, treating the ensembled prediction as a pseudo label . This pseudo label is then used to filter out responses with conflicting predictions, ensuring consistency. Following the best practices outlined in §3.3, we apply PRM as a reranker to select the top-2 responses among those with the predicted answer . These unlabeled prompts are then mixed with labeled prompts for self-evolving training. Additionally, since learning from unlabeled prompts is more challenging for policy models, we investigate the optimal stage to introduce them into training to better understand their impact on model performance. We maintain a training interval of 45k prompts and adjust when unlabeled prompts are introduced into the self-evolving training process. Specifically, we introduce unlabeled prompts after [0%, 25%, 50%, 75%] of the total training process.
A Glimpse at Unlabeled Prompts: Potential Efforts to Make Them Effective
Table 3 presents the results of incorporating unlabeled prompts with and without oracle reward signals. When training relies solely on oracle reward signals without integrating the PRM, continuous self-evolving with unlabeled prompts outperforms standard continuous self-evolving trained only on labeled prompts in the out-of-domain test but underperforms in the in-domain test. This indicates that additional prompts help the model generalize better to underrepresented questions but also increase the risk of forgetting previously learned information.
@@ A Glimpse at Unlabeled Prompts: Potential Efforts to Make Them Effective
However, after combining with our PRM, all policy models perform worse than our best model trained exclusively on labeled prompts in both benchmarks, even when oracle reward signals are provided. Based on the analysis in §3.3, this occurs since our PRM is unable to verify responses without ground-truth answers, and its generalization remains a concern.
When examining the timing for introducing unlabeled prompts, we find that adding them from the beginning helps mitigate the negative impact on model performance, compared to introducing them midway through the process. However, when unlabeled prompts are introduced later in the training process, they participate less in the overall training, leading to better results simply due to their limited involvement. This suggests that, without sufficient surrogate supervision (e.g., reward signals), introducing unlabeled prompts during the middle stages of self-evolving training can harm the process, potentially causing a deviation in the policy model’s distribution.
4 Dynamics of Self-Evolution and the Final Recipe

### v3

@@ Abstract
Self-evolving training—where models iteratively learn from their own outputs—has emerged as a key approach for complex reasoning tasks, addressing the scarcity of high-quality chain-of-thought data. However, its effectiveness in multimodal reasoning, a domain more intricate than text-only reasoning, remains underexplored, and the understanding of critical factors in this training paradigm remains limited. Furthermore, a central challenge for this training method is performance saturation, which impedes further improvements and scalability. Inspired by reinforcement learning (RL), in this paper, we reframe self-evolving training for multimodal reasoning through the lens of RL, identifying three pivotal factors: Training Method, Reward Model, and Prompt Variation. Through systematic analysis, we establish relatively optimal design principles that significantly enhance multimodal reasoning capabilities. Moreover, delving deeper into training dynamics, we uncover the roots of saturation and propose a new automatic balancing mechanism to mitigate this limitation. Building on these insights, we propose M-STaR (Multimodal Self-evolving Training for Reasoning), a framework that achieves consistent performance gains across models of varying sizes and diverse benchmarks. All resources are made publicly available at https://mstar-lmm.github.io.
@@ 1 Introduction
Multimodal reasoning is a fundamental skill in many real-world applications, such as intelligent agents (Liu et al., 2024c), robotics (Li et al., 2023; Liu et al., 2024b), and autonomous driving (Yang et al., 2023). It requires Large Multimodal Models (LMMs) to understand various modalities beyond text. For example, visual mathematical reasoning (Lu et al., 2023) challenges models to analyze complex figures, diagrams, and charts, leveraging the provided information to perform reasoning tasks.
@@
Despite the critical role of mutlimodal reasoning, the availability of human-annotated thought processes in multimodal scenarios remains limited, hindering the learning of multimodal reasoning. Consequently, self-evolving training, which utilizes model’s own generation ability to iteratively tune and improve itself without external annotated data, has emerged as an appealing candidate to facilitate reasoning abilities. While research on self-evolving training has primarily focused on the text-only settings (Hosseini et al., 2024; Sun et al., 2024; Shao et al., 2024), its application in the multimodal domain, especially for reasoning tasks, has been limited with only a few sporadic examples (Fang et al., 2024; Dubey et al., 2024; Deng et al., 2024), and a unified framework has yet to be established.
@@
Inspired by reinforcement learning (RL), in this paper, we reframe self-evolving training through the lens of RL, identifying three factors that are critical inside self-evolving training: the training method, the use of reward model, and prompt variation. Through massive controlled studies, we (1) propose a continuous self-evolving training scheme to reduce the gap towards full online learning and outperforms other iterative baselines (§3.2); (2) train the first multimodal, process-based reward model for multimodal reasoning and demonstrate its usefulness in further enhancing performance (§3.3); and (3) find that adding more unlabeled queries helps only when having perfect reward signals (e.g., the oracle groundtruth answers), and it hurts the performance if the reward model does not generalize well on unseen data (§3.4). Beyond static design principles, we investigate the training dynamics, revealing how performance saturation stems from diminishing exploration potential during training. To address this, we introduce a new metric that bridges exploration and exploitation, and propose an automatic balancing mechanism that dynamically adjusts the sampling temperature to sustain exploration-exploitation trade-offs.
@@
Combining all the recipes concluded through separate, controlled studies, we propose our self-evolving training algorithm named as M-STaR (Multimodal Self-evolving Training for Reasoning). Our experimental results on 5 different multimodal reasoning benchmarks, including MathVista, M3CoT, MMStar, MMBench and AI2D, show that this strategy, which incorporates both optimized static design choices and dynamic adjustments, effectively mitigates exploration loss during training and enhances performance universally for models with varied sizes such as MiniCPM-V-2.5 (8B), Phi-3.5-Vision (4B) and InternVL2 (2B).
@@ 2 Overview of Self-Evolving Training for Multimodal Reasoning
where the ground-truth answer can be empty, for example, when dealing with unlabeled inputs, the reward model must be able to score independently.
The Design Spaces Through the lens of Eq. 2, we can identify three dominant factors that influence the training process, (1) training method: the training algorithms to perform this iterative process vary as well. For example, while Gulcehre et al. (2023); Xu et al. (2024b) initialize the model from the last checkpoint at each iteration, Zelikman et al. (2022); Singh et al. (2023) argue that initializing from the beginning checkpoint reduces overfitting and gives better performance empirically. (2) reward model: the design of reward function . (3) prompt variation: whether to incorporate additional unlabeled inputs without into training. Next, we investigate these three design spaces, aiming to summarize the best practices for each factor.
@@ 3.1 General Setup
Models We base our main exploration on MiniCPM-V-2.5 (8B) (Yao et al., 2024), and we also validate the final design choice for each component on two extra models with different sizes: Phi-3.5-Vision (4B) (Abdin et al., 2024) and InternVL-2 (2B) (Chen et al., 2024c). The details of these models can be found in Appendix A. To make the analysis process easier to understand, we mainly present the results of MiniCPM-V-2.5 in this section, while we include the results of the other models in §4.2.
Datasets We utilize MathV360K (Shi et al., 2024), a high-quality and diverse multimodal reasoning dataset as our seed training dataset. Specifically, we downsample half of the examples (180K) from it to serve as our labeled training set, while setting aside the remaining half as a unlabeled training set by not using the answers in it. For evaluation, we split 750 samples from the unlabeled part of MathV360K as the in-domain (ID) testset. For our out-of-domain (OOD) testset we use the testmini split of MathVista (Lu et al., 2023), a comprehensive benchmark encompassing a wide range of multimodal reasoning tasks, including visual question answering, figure-based question answering, science question answering, and more. We also keep an non-overlapping 250 samples from MathV360K as the global validation set in training.
Main Training Settings Before self-evolving training, a self-warmup stage (Appendix B) is applied to enhance the CoT generation ability of all models as many LLMs tend to output final answers directly without CoT. We adopt most of the training settings from Yao et al. (2024) (see Appendix C), using a constant learning rate of and training for 10K steps across all experiments. During all rollout phases in training, we sample 16 responses per query and set the sampling temperature to 1.0. Unless explicitly stated otherwise, we follow existing practices (Singh et al., 2023; Zelikman et al., 2022) and only use the labeled training data.
@@ 3.2 Training Methods
Setup We perform controlled experiments to study the effect of different training methods, thus in this experiment we use the labeled dataset only and simply adopt the binary exact-match reward between ground-truth answer and the generated answer. We compare with the most common iterative self-evolving algorithms ReSTEM (Singh et al., 2023) and iterative RFT, which are specific instantiations of our training methods design space. To study the effect of iteration interval in the proposed continuous self-evolving, we experiment with different percentage of all the queries per iteration, varying from [6.25%, 12.5%, 25%, 50%, 100%].
| Method | Interval (%) | MathV360K | MathVista | ||
@@
| Iterative RFT | 100% | 42.3 | 55.7 | ||
| RestEM | 100% | 42.3 | 55.1 | ||
| Continous Self-Evolving | 100% | 42.2 | 56.7 | ||
| 50% | 43.1 | 56.2 | |||
| 25% | 43.1 | 57.2 | |||
| 12.5% | 42.3 | 56.1 | |||
| 6.25% | 41.0 | 56.8 |
Results Table 1 presents the experimental results of various training methods. Overall, initializing training from the last policy model checkpoint and maintaining a continuous optimization process contribute most significantly to the effectiveness of self-evolving training, particularly on MathVista. Continuous self-evolving achieves the best performance both on the in-domain MathV360K test set, with 43.1%, and on the OOD test set, MathVista, with 57.2%. We also see the importance of maintaining a proper interval to traverse the data queries. With a large interval, the training method becomes closer to an offline one, and the model cannot get timely updates on data matching its current output distribution. On the other hand, switching over the Improve and Generate steps too frequently makes the learning process unstable, leading to a lower score, especially on the in-domain test set. The strategy of continuous self-evolving with proper intervals also works for other smaller models, as shown in Table 6 compared with representative baselines, indicating its effectiveness and generalizability across different model sizes.
@@ 3.3 Reward Models
Setup We conduct controlled experiments to assess the impact of incorporating the Process Reward Model (PRM) into self-evolving training and explore how best to utilize it. Notably, before applying PRM, responses are pre-filtered based on their final answers to ensure consistency and quality during training. To train our PRM, we use Monte Carlo rollouts starting from prefixes with partial reasoning steps (Wang et al., 2024) to generate the training data. Specifically, we sample 16 responses per question and complete each step 8 times to obtain step-level annotations and more details can be found in Appendix D. We evaluate two different operations: (1) Top-K: Pick the top-K correct responses according to their reward scores, and (2) Filtering by a Threshold : Filtering out sampled responses with lower aggregated rewards than . The optimal value of is set by grid searching on the validation set. Additionally, we investigate how varying the value of in affects training, as it represents a trade-off between the quality and diversity of the samples. According to §3.2, we fix training methods as continuous self-evolving with 45k interval and set continuous self-evolving, with or without randomly selected correct responses as our baselines.
Results Table 2 presents the results of integrating the PRM into self-evolving training, along with the impact of different choices. Continuous Self-Evolving with PRM using achieves the best performance in both the ID and OOD tests, with scores of 45.3% and 59.2%, respectively. Compared to training without PRM, most instances of self-evolving training with PRM show improved performance, especially in the OOD test. Interestingly, randomly selecting a subset of correct responses actually leads to worse performance than continuous self-evolving, suggesting that even correct answers can be noisy. Random selection may increase the proportion of these noisy samples, undermining the effectiveness of self-evolving training.
In terms of leveraging PRM, we found that using to select the a fixed number of best responses with high-quality intermediate steps outperforms threshold-based filtering. The results also highlight the importance of balancing the quality and diversity of sampled responses. Selecting strikes this balance well, ensuring both response diversity and high-quality reasoning steps for each question. Similar to the results in §3.2, we also see improvement when involving PRM on smaller models in Table 6, Appendix H.
@@
| Cont. Self-Evolving | - | 43.1 | 57.2 | |
| + Random | 41.0 | 55.5 | ||
@@
What makes PRM work for self-evolving training? To pursue deeper insights into the role of PRM in self-evolving training, we conduct an analysis presented in Figure 2. Based on the results from §3.3, we explore PRM’s impact from two key perspectives: (1) Can PRM help the model to select out correct responses among multiple rollouts? (2) How different are the Top 2 and the rest correct solutions re-ranked by reward scores? We use the first checkpoint after warmup as policy model to sample 16 responses for each question in the validation set with temperature= and reveal the behaviors of PRM in these samples.
@@
To understand why our PRM can still significantly contributes to self-evolving training despite its weaker verification abilities, we analyzed the distribution of other metrics for the top-2 selected responses compared to other correct responses. We approached this from two perspectives: the average number of reasoning steps, and how much a response is directly relevant to the question (see Appendix E), since we do not find incorrect steps but find some irrelevant steps after randomly checking some examples. We found the responses re-ranked by our PRM generally have fewer reasoning steps (Figure 6 in Appendix F) and more relevant to the query (Figure 2(b)). This highlights the precision of our PRM in recognizing genuinely high-quality responses. Therefore, our PRM acts as an effective reranker to identify top-quality responses. This is especially critical when responses are already filtered by ground-truth answers, and the ability to accurately assess the quality of reasoning steps beyond accuracy becomes vital.
@@ 3.4 Prompt Variation
Skylines: Unlabeled Prompts with Oracle Reward Signals The coupling of these additional factors introduces complexity, making the effective use of unlabeled prompts less predictable. Therefore we start by establishing a baseline with “skyline” experiments, where both the unlabeled prompts and their ground truth answers are available but not used during the SFT phase. These unlabeled prompts with oracle reward signals serve as an intermediate difficulty between fully unlabeled and labeled prompts, providing insight into the challenges of training with unlabeled data.
Unlabeled Prompts We incorporate unlabeled prompts into self-evolving training. To ensure the quality of sampled responses for these prompts, we use weighted voting to ensemble the predictions from different responses, treating the ensembled prediction as a pseudo label . This pseudo label is then used to filter out responses with conflicting predictions, ensuring consistency. Following the best practices outlined in §3.3, we apply PRM as a reranker to select the top-2 responses with the predicted answer . These unlabeled prompts are then mixed with labeled ones for self-evolving training. Additionally, since learning from unlabeled prompts is more challenging for policy models, we investigate the optimal timing to introduce them into training as well. We maintain an interval of 45k prompts and adjust when unlabeled prompts are introduced into the training process. Specifically, we introduce unlabeled prompts after [0%, 25%, 50%, 75%] of the total training process.
A Glimpse at Unlabeled Prompts: Potential Efforts to Make Them Effective Table 3 presents the results of incorporating unlabeled prompts with and without oracle reward signals. When training relies solely on oracle reward signals without the PRM, continuous self-evolving with unlabeled prompts outperforms standard continuous self-evolving trained only on labeled prompts in the out-of-domain test but underperforms in the in-domain test. This indicates that additional prompts help the model generalize better to underrepresented questions but also increase the risk of forgetting previously learned information. However, after combining with our PRM, all policy models perform worse than our best model trained exclusively on labeled prompts in both benchmarks, even when oracle reward signals are provided. Based on the analysis in §3.3, this occurs since our PRM is unable to verify responses without ground-truth answers, and its generalization remains a concern.
@@
When examining the timing for introducing unlabeled prompts, we find that adding them from the beginning helps mitigate the negative impact on model performance, compared to introducing them in midway. However, when unlabeled prompts are introduced later in the training process, they participate less in the overall training, leading to better results simply due to their limited involvement. This suggests that, without sufficient surrogate supervision (e.g., precise reward signals), introducing unlabeled prompts into self-evolving training can harm the process, potentially causing a deviation in the policy model’s distribution.
4 Dynamics of Self-Evolution & Final Recipe